OpenStudy (anonymous):

Given y-2xy=x^3, determine the equation of the tangent line where the point of tangency occurs at x=3

OpenStudy (anonymous):

I'll be right back tomorrow. I have to go to sleep. If someone can help me start the question that will be right. Appreciated it! :)

OpenStudy (rational):

You need a `point` on the line and `slope` to write the equation of tangent line

OpenStudy (rational):

derivative at x=3 gives the slope of tangent line at x=3

OpenStudy (rational):

y-2xy = x^3 y(1-2x) = x^3 y = x^3/(1-2x) dy/dx = ?

OpenStudy (anonymous):

@ganeshie8 i need help to figure out dy/dx

ganeshie8 (ganeshie8):

You have : \[y = \dfrac{x^3}{1-2x}\] maybe use quotient rule to find \(\dfrac{dy}{dx}\)

OpenStudy (anonymous):

okay, let me try it

OpenStudy (anonymous):

\[y'=\frac{ -4x^3+3x^2 }{ (1-2x)^2 }\] is that correct?

ganeshie8 (ganeshie8):

Yes! plugin x=3 since you want the slope of tangent line at x=3

OpenStudy (anonymous):

is it -27/5?

OpenStudy (anonymous):

@ganeshie8 how do i figure the equation?

ganeshie8 (ganeshie8):

doesn't look correct, check again

ganeshie8 (ganeshie8):

wolfram says -81/25 http://www.wolframalpha.com/input/?i=%28x%5E3%2F%281-2x%29%29%27+at+x%3D3

OpenStudy (anonymous):

oh i got -81/25 now @ganeshie8

ganeshie8 (ganeshie8):

good so you know the "slope" of line you just need a point on the line to write the equation, yes ?

OpenStudy (anonymous):

yes!

ganeshie8 (ganeshie8):

plugin x=3 in to the given equation and solve y

ganeshie8 (ganeshie8):

y-2xy=x^3 plugin x=3

OpenStudy (anonymous):

y= -27/5?

ganeshie8 (ganeshie8):

Yep so you have : slope of line = -81/25 point on line = (3, -27/5)

OpenStudy (anonymous):

oh so it will -81/25=(y+27/5)/(x-3)

ganeshie8 (ganeshie8):

Looks good!

OpenStudy (anonymous):

so it would be 0=81x-25y-108?

OpenStudy (anonymous):

thanks! @ganeshie8